Limx→∞ -(x+1)(e1/(x+1) -1)


1
Expert's answer
2021-06-07T03:33:11-0400

We need to calculate "lim(x\to\infty) -(x+1)(e^\frac {1} {(x+1)} -1)"


"=lim(x\to\infty)(-1)\frac {e^\frac {1} {(x+1)} -1 } {\frac {1} {(x+1)}}"

It is in "\frac 0 0" form after putting "\infty" in place of "x" .So L hospital's rule is applicable here.


="(-1) lim(x\to\infty)\frac {\frac {d} {dx}(e^\frac {1} {(x+1)} -1)} {\frac {d} {dx}\frac {1} {(x+1)}}"


"=(-1)lim(x\to\infty)\frac {e^\frac {1} {(x+1)}. \frac {-1} {(x+1)^2}} {\frac {-1} {(x+1)^2}}" "(\because \frac {d} {dx} \frac u v" "=\frac {(v.u'-u.v')} {v^2} )" So ("\frac {d} {dx} \frac {1} {x+1} =\frac {-1} {(x+1)^2})"

( "\frac {d} {dx} e^x=e^x)"



"=(-1)lim(x\to\infty)e^\frac {1} {(x+1)}"


"=(-1) e^\frac {1} {\infty}"


"=(-1)e^0"


"=(-1).1 =-1" (Ans)